# Superlatives in Context: Explicit and Implicit Domain Restrictions for Superlative Frames

### v1

@@ 5.4 Ambiguous Superlatives
The main ambiguity present for superlatives is the distinction between absolute versus relative interpretations (see Sec. 2.3). In an effort to analyze how sensitive our model is to discourse cues which could help to disambiguate between different readings, we perform the following experiment. We manually curated 20 sentences in which the superlative comparison is ambiguous. Many of these sentences are (synthetic) example sentences found in formal semantic literature. Additionally, for each of the 20 instances, we added context that strengthens a certain reading, such as absolute vs. relative readings. As follows we show an example of our synthetic test set.
  1. 1.
    John put the tallest plant on the table.
  2. 2.
@@
The first sentence is ambiguous in that it could be read as either absolute or relative, i.e. restricted by the putting event or not. Given the additional context 1, the relative reading is strengthened: PUT(e, AGENT=Tom & John & Mary, PATIENT=plants, DESTINATION=table). We run the T5-3b model trained to predict the CS slot on the synthetic test data.
@@ 5.5.1 Probability Given Context
  1. 1.
    John is angriest at Mary.
  2. 2.
@@ 5.6 Entropy
The instances with the highest difference in entropy between the no-context and with-context setups, reveal some patterns. Cases where entropy is low for with-context but high for no-context tend to be extremely underspecified superlatives, usually in dialogue turns, where a sentence might simply say “Which one is best?". The eventive and property CS predictions have higher entropy for the full-context model, as it more frequently predicts eventive interpretations in its beam, compared to the no-context model.
@@ 6 Related Work
Multiple works have studied ambiguity, but none of them have specifically focused on superlatives: Cui et al. (2022) look at generalized quantifier ambiguity in multilingual NLI data, Liu et al. (2023) look at sentence ambiguity and its effect on entailment relations and Stengel-Eskin et al. (2023) introduce a framework to translate ambiguous statements to formal representations.
@@ 7 Conclusion
Superlative comparisons are interesting because their interpretation is closely tied to the context they appear in and because many of their components are often implicit. This paper provides a comprehensive study of superlatives, by proposing a new, unified annotation scheme and an annotated superlatives dataset, SuperSem. We further perform a set of experiments which analyze how models interpret superlative comparisons and how they are able to incorporate context restrictions. And with help of the SuperSem dataset we could investigate how ambiguity and context interact, for superlative interpretations.

### v2

@@ 5.4 Ambiguous Superlatives
[…]
@@
The first sentence is ambiguous in that it could be read as either absolute or relative, i.e. restricted by the putting event or not. Given the additional context 1, the relative reading is strengthened: PUT(e, AGENT=Tom & John & Mary, PATIENT=plants, DESTINATION=table).
We run the T5-3b model trained to predict the CS slot on the synthetic test data.
@@ 5.5.1 Probability Given Context
[…]
. John is angriest at Mary.
@@ 5.6 Entropy
The instances with the highest difference in entropy between the no-context and with-context setups, reveal some patterns. Cases where entropy is low for with-context but high for no-context tend to be extremely underspecified superlatives, usually in dialogue turns, where a sentence might simply say “Which one is best?”. The eventive and property CS predictions have higher entropy for the full-context model, as it more frequently predicts eventive interpretations in its beam, compared to the no-context model.
@@ 6 Related Work
Multiple works have studied ambiguity, but none of them have specifically focused on superlatives: Cui et al. (2022) look at generalized quantifier ambiguity in multilingual NLI data, Liu et al. (2023) look at sentence ambiguity and its effect on entailment relations, and Stengel-Eskin et al. (2023) introduce a framework to translate ambiguous statements to formal representations.
@@ 7 Conclusion
Superlative comparisons are interesting because their interpretation is closely tied to the context they appear in and because many of their components are often implicit. This paper provides a comprehensive study of superlatives, by proposing a new, unified annotation scheme and an annotated superlative dataset, SuperSem. We further perform a set of experiments which analyze how models interpret superlative comparisons and how they are able to incorporate context restrictions. And with the help of the SuperSem dataset we could investigate how ambiguity and context interact for superlative interpretations.